Find the solutions to the equation: x-5/x=4
Answer by checkley71(8403)

X-5/X=4 OR (X^2-5)/X=4 OR X^2-5=4X OR X^2-4X-5=0 OR(X-5)(X+1)=0 OR X=5 & X=-1